Calculate the amount of heat released (kJ) in the complete oxidation of 55.8 g of Fe metal at standard condition.

4 Fe(s) + 3 O2(g) → 2 Fe2O3(s) ΔH° = -1648 kJ
1
Expert's answer
2019-09-30T05:13:50-0400

4 Fe(s) + 3 O2(g) → 2 Fe2O3(s) ΔH° = -1648 kJ

n=1 mol(55.8/55.8)

4 -------1648 kJ

1--------x

x= 1648/4= 412 kJ.